I am having trouble writing a algebra problem for the following:
Sue has 15,000,000 and gives Fred 200,000. Fred puts it in a bank at 5.9% interest. Sue puts her money in the freezer and pulls out 50,000 a month Sue spends 600,000 a year. After how many years would Sue and Fred have the same amount of money? Could you please help with an algebraic equation.

Answer by stanbon(75887)

Sue has 15,000,000 and gives Fred 200,000. Fred puts it in a bank at 5.9% interest. Sue puts her money in the freezer and pulls out 50,000 a month Sue spends 600,000 a year. After how many years would Sue and Fred have the same amount of money? Could you please help with an algebraic equation.
-------------
Let "t" be # of years
-------------
Value of Fred's money: F(t) = 200,000*0.059*t
Value of Sue's money: S(t) = 14,800,000-600,000t
-----------------------
EQUATION:
F(t) = S(t)
200,000*0.059t = 14,800,000-600,000t
11800t = 14,800,000-500,000t
511,800t = 14,800,000
t = 28.92 years
========================
Cheers,
Stan H.